OpenStudy (anonymous):

g(x) = 7x - 2 h(x) = 4x2 + 7 Find: (g + h)(-3)

OpenStudy (anonymous):

g(-3) +h(-3) = 7*(-3) -2 + 4*(-3)^2 +7 = 20